# thanks: Corresponding author

### v1

@@ V Operator delocalization through Krylov complexity
The analogy of a particle hopping on the Krylov chain can be made more direct by examining the probability distribution of the particle’s position, given by the Krylov wavefunction, $|\phi_{n}|^{2}$. This is reported in Fig. 10 for multiple $\tilde{M}$ at a late time of $t\approx 2\times 10^{5}$. A significant weight in $|\phi_{n}|^{2}$ at large $n$ indicates that the time-evolved operator, $\mathcal{O}(t)$, has spread to more complex operators in the Krylov basis. We observe that $|\phi_{n}|^{2}$ decays rapidly with $n$ at large $\tilde{M}$, indicating that the operator remains localized near the beginning of the Krylov chain. In contrast, it shows a much slower decay in the chaotic regime, signifying delocalization across the chain. The qualitative trend holds for all operator sizes $|\mathcal{O}|$. Thus, the single-particle picture provides a consistent and intuitive explanation for the operator complexity. The transition to chaos is reflected in the eventual delocalization of the operator across the entire Krylov chain.
VI Anomalous relaxation of initial product states
We now turn to the dynamical relaxation of quantum states under unitary time evolution. In closed quantum many-body systems, a subsystem’s approach to equilibrium is driven by the scrambling of local information, a process whose efficiency was characterized in the previous section via Krylov complexity. In this section, we focus on how certain simple initial states relax under this unitary dynamics. An intuitive expectation for individual states is that the time to reach equilibrium should monotonically increase with the initial state’s “distance” from the equilibrium ensemble. In a classical context, a “hotter” system is farther from ambient temperature compared to a “cooler” system. So, we can expect that a “cooler” state, being “closer” to equilibrium, would relax faster than a “hotter” one. However, it has been found that certain initial conditions can defy this intuition, giving rise to anomalous thermalization behaviors. Broadly, this phenomenon is studied as the Mpemba effect, first reported in water [75], where initially hotter systems freeze faster than cooler ones under a quench.
To investigate the QME in this random graph system, we first establish a quench protocol and a metric to quantify the thermalization. Unlike studies that track the restoration of a global symmetry, our Hamiltonian (Eq. 1) with $h\neq 0$ lacks such a conserved quantity. We therefore adopt a more direct approach [85] to quantify relaxation. Our metric is the trace distance, $\Delta^{(1)}$ (Eq. 3), between the time-evolved reduced density matrix and its corresponding diagonal ensemble. This diagonal ensemble represents the infinite-time equilibrium ensemble ($\rho_{A}(\infty)$) given as,
where $H=\mathcal{U}\tilde{E}U^{\dagger}$ for Hamiltonian $H$ and $\tilde{E}$ is a diagonal matrix containing the eigenvalues of $H$. $B$ indicates the complement of subsystem $A$. To simplify the parameter space of the initial state, we prepare a family of translation-invariant product states $\psi(\theta,\phi)$ in real space as,
$E(\theta,\phi)$ depends only on the coupling parameters of the Hamiltonian but not on the connectance (This is due to the normalization factors in the denominator of Eq. 1). The distribution of initial state energy as a function of $(\theta,\phi)$ is shown in Fig. 11. The red and blue region shows high and low energy states, respectively, with a crossover region near 0 energy where states correspond to infinite temperature. Consider two initial states $\psi(\theta_{1},\phi_{1})$ and $\psi(\theta_{2},\phi_{2})$ such that $\Delta^{(1)}_{\theta_{1},\phi_{1}}(t=0)>\Delta^{(1)}_{\theta_{2},\phi_{2}}(t=0)$ where trace distance is always taken from the respective diagonal ensembles. Existence of QME implies for $t>t_{M}$, $\Delta^{(1)}_{\theta_{1},\phi_{1}}(t)<\Delta^{(1)}_{\theta_{2},\phi_{2}}(t)$, where $t_{M}$ is referred to Mpemba time. Here, $\Delta^{(1)}_{\theta_{1},\phi_{1}}(t=0)$ is the initial “hotter” state due to the initial higher trace distance of the subsystem reduced density matrix from the corresponding equilibrium ensemble.
Our analysis focuses on two distinct classes of initial states(Eq. 18), parametrized by $\theta$ and marked in Fig. 11. The first class, with $\theta=0.5\pi$, spans a wide range of initial energies as $\phi$ is varied. In contrast, the second class, with $\theta=0.95\pi$, consists of nearly-polarized states whose initial energies are largely independent of $\phi$.
Figure 12: Anomalous relaxation in the chaotic regime at $\tilde{M}=0.4$ for a system with $L=12$. (a) The Inverse Participation Ratio (IPR) as a function of the initial state parameters $(\theta,\phi)$. The red and blue vertical lines at $\theta=0.5\pi$ and $\theta=0.95\pi$ indicate the two classes of states studied. (b) Relaxation dynamics, measured by the trace distance $\Delta^{(1)}(t)$ from the diagonal ensemble, for the nearly-polarized states ($\theta=0.95\pi$). (c) The distribution of overlaps, $|\langle E_{n}|\psi\rangle|^{2}$, with energy eigenstates for several initial states from the $\theta=0.5\pi$ class. The inset shows the corresponding IPR for each state. (d) Relaxation dynamics for the $\theta=0.5\pi$ class, showing a clear crossing indicative of the QME. For all plots, the subsystem size is $L_{A}=4$, and all data are averaged over $500$ graph realizations using the Hamiltonian in Eq. 1 with $g=1.0$ and $h=0.1$.
Starting with these two classes of initial states (identified by $\theta$), we time evolve with a specific graph realization of Hamiltonian in Eq. 1 with $h=0.1,g=1$. We first consider $\tilde{M}=0.4$, where the model is expected to be chaotic. Figure 12(b) and (d) shows time evolution of $\overline{\Delta^{(1)}}$ for $\theta=0.5\pi$ and $\theta=0.95\pi$, respectively. All data are presented after averaging over $500$ graph realizations starting with the same set of initial states. We observe a clear crossing in $\Delta^{(1)}$ in Fig. 12(d) where the initial states are taken such that $\theta=0.5\pi$. Such crossing is absent in Fig. 12(b) where $\theta=0.95\pi$ and relaxation to equilibrium looks almost independent of the initial state at the late time.
where $c_{n}=\langle\psi|E_{n}\rangle$. At $t=0$, $\Delta^{(1)}(0)=\frac{1}{2}\left\|\sum_{n\neq m}c_{n}c_{m}^{*}\mathrm{Tr}_{\mathrm{B}}(|E_{n}\rangle\langle E_{m}|)\right\|_{1}$ i.e. the trace distance contains $O(d^{2})$ terms weighted by the overlaps of initial state with eigenstates. As was pointed out in Ref [85], the relaxation rate of states can be related to the number of nonzero coefficients $c_{n}$, quantified by the inverse participation ratio (IPR) given as,
$\mathrm{IPR}(\psi)=\sum_{i=1}^{d}|\langle E_{i}|\psi\rangle|^{4}.$ (21)
$|\Delta^{(1)}(t)|^{2}$ is the sum of squared norms of the matrix elements of $(\rho_{A}-\rho_{A}{(\infty}))$. For an initial state $\psi$ which has overlap with a large number of energy eigenstates (i.e., smaller IPR), there are large number of nonzero $c_{n}$s. Each $c_{n}$ is smaller due to normalization of the initial state and, at the same time, there are a larger number of fluctuating phases in the expansion of each matrix elements of $(\rho_{A}-\rho_{A}{(\infty}))$. As a result, the temporal fluctuation in trace distance as the system relaxes are weaker in the case where IPR is smaller. For states with high IPR, a larger number of $c_{n}$s will be close to 0, which results in temporal oscillation in trace distance from equilibrium. This causes a slower equilibration.
In Fig. 12, for the “cooler” initial states, the initial distance from the diagonal ensemble is lower as compared to “hotter” states. Though, we can not prove this, we attempt to provide a plausibility argument. First, we consider the low IPR states. This states have finite expectation value for Pauli observables. The diagonal ensemble constructed with this initial state will be a mixture of highly entangled mid-spectrum eigenstates. However, the diagonal ensemble will have 0 expectation value of such observables (as it is an almost uniform mixture of mid-spectrum eigenstates, which resemble infinite temperature states according to ETH). This suggests that the diagonal ensemble is not a good approximation at the initial time for the direct product state, resulting in a large initial trace distance. Note that, states close to the ground state are being approximated by a diagonal ensemble which, consists of low-lying eigenstates that can have finite local expectation values.
Figure 13: Anomalous relaxation in the near-integrable regime at $\tilde{M}=0.8$ for a system with $L=12$. (a) The IPR as a function of the initial state parameters $(\theta,\phi)$. The red and blue vertical lines indicate the two classes of states studied. (b) Relaxation dynamics, $\Delta^{(1)}(t)$, for the nearly-polarized states ($\theta=0.95\pi$). (c) The distribution of overlaps, $|\langle E_{n}|\psi\rangle|^{2}$, with energy eigenstates for several initial states from the $\theta=0.5\pi$ class. (d) Relaxation dynamics for the $\theta=0.5\pi$ class, showing persistent oscillations and multiple crossings. The inset provides a zoomed-in view of the initial relaxation. For all plots, the subsystem size is $L_{A}=4$, and all data are averaged over 500 graph realizations using the Hamiltonian in Eq. 1 with $g=1.0$ and $h=0.1$.
The clear observation of the QME in the chaotic regime is driven by the broad delocalization of initial states across the energy spectrum. A natural question is how this phenomenon behaves as the system approaches the integrable limit at high connectance ($\tilde{M}\to 1$), where the eigenspectrum deviates from chaotic WD statistics. We investigate this interplay at $\tilde{M}=0.8$ (Fig. 13). While the correlation between IPR and energy persists (Fig. 11 and 13(a) ), the overall range of IPR values is smaller, indicating that the initial states are more localized in the eigenbasis as $\tilde{M}\to 1$. Consequently, the nearly-polarized $\theta=0.95\pi$ states still show no strong signature of the QME (Fig. 13(b)). For the $\theta=0.5\pi$ states, however, the dynamics are markedly different from the chaotic case. Instead of a smooth decay, the trace distance exhibits persistent late-time oscillations, as seen in Fig. 13(d). We expect this to be caused by the emergence of approximately conserved charges close to an integrable point. In such near-integrable models, the QME is understood to arise from the coherent propagation of stable quasiparticles, which leads to persistent oscillations and multiple crossings of the relaxation curves [82, 87]. Our results show these characteristic signatures of multiple crossings and oscillatory behavior, and we observe the QME within the timescale studied (Fig. 13(d)).
VII Conclusion and Outlook
@@
In this work, we have investigated the emergence of quantum chaos and complexity in the mixed-field Ising model defined on random Erdős–Rényi graphs, with a particular focus on how the connectance parameter $\tilde{M}$ controls the dynamical and spectral properties of the system. Using a combination of diagnostics, including projected ensemble (PE), partial spectral form factor (pSFF), and Krylov complexity (KC), we provide a detailed characterization of the localized to chaotic to integrable crossover driven by changes in connectance.
The slowdown of deep thermalization at the extreme limits of connectance is signaled by the distance of the PE from the Haar ensemble. In the intermediate chaotic regime, the pSFF exhibits the characteristic ramp-and-plateau behavior indicative of RMT statistics, while the Krylov complexity correspondingly reveals enhanced operator spreading and higher late-time saturation values. Although the studies presented here are limited to modest system sizes, our suite of chaos indicators provides clear and consistent signatures of the crossover from localization to chaos.
We also investigate the relaxation dynamics of initial product states, where initial states further from their equilibrium (“hotter”) thermalize faster than the initial states closer to their equilibrium (“cooler”) under unitary evolution. As $\tilde{M}\to 1$, the signature of the effect is modified, with persistent oscillations causing multiple crossings in the trace distance. Our analysis has focused on the ferromagnetic Ising model. While the dynamical signatures of chaos are expected to be independent of the sign of the couplings, the relaxation dynamics, particularly for specific low-energy initial states, may be altered in the anti-ferromagnetic case, which is a potential direction for future study.
Our results are directly relevant for benchmarking and characterizing thermalization in near-term quantum annealing and simulation platforms. Furthermore, a useful avenue for future work is to focus on the dynamics within the low-energy sector of the spectrum. Since the performance of quantum annealing is sensitive to the localization of low-lying eigenstates [26], applying our suite of chaos diagnostics to this spectral edge could reveal a more direct link between chaos and algorithmic performance.
@@ Acknowledgements.
Appendix A Lanczos algorithm

### v2

@@ V Operator delocalization through Krylov complexity
The analogy of a particle hopping on the Krylov chain can be made more direct by examining the probability distribution of the particle’s position, given by the Krylov wavefunction, $|\phi_{n}|^{2}$. This is reported in Fig. 13 for multiple $\tilde{M}$ at a late-time of $t\approx 2\times 10^{5}$. Finite $|\phi_{n}|^{2}$ even at large $n$ indicates that $\mathcal{O}(t)$ involves high-complexity operators in the Krylov basis. We observe that $|\phi_{n}|^{2}$ decays rapidly with $n$ at large $\tilde{M}$, indicating that the operator remains localized near the beginning of the Krylov chain. In contrast, it shows a much slower decay with $n$ in the chaotic regime, signifying delocalization across the chain. The qualitative trend holds for all operator sizes $|\mathcal{O}|$.
Experimental access to late-time Krylov complexity is challenging, and the development of scalable experimental protocols is still in the nascent stage [29, 77]. Thus even though asymptotic values of Krylov complexity shows clear signatures of chaos as seen in Fig. 12 experimentally reproducing them or obtaining a clear scaling is difficult. However early time behavior of Krylov complexity can be captured in current devices by measuring two-point correlator $F(t)=\langle\mathcal{O}(t)\mathcal{O}(0)\rangle$. Since $F(t)$ is an even function of time (due to cyclicity of trace), it can be expanded in even powers of $t$ as
$F(t)=\sum_{n=0}^{\infty}\frac{(-1)^{n}t^{2n}}{(2n)!}\mu_{2n},$ (25)
where the ‘moments’ $\mu_{2n}$ are related to the adjoint action of the Liouvillian on the operators as $\frac{\rm{Tr}(\mathcal{O}\mathcal{L}^{2n}\mathcal{O})}{\rm{Tr}(\mathcal{O}^{2})}$ are the moments of the correlation function. The Lanczos coefficients $\{b_{n}\}$, which govern the growth of the operator in Krylov space, are entirely determined by these moments via a Gram-Schmidt orthogonalization of the sequence $\{\mathcal{O},\mathcal{L}\mathcal{O},\mathcal{L}^{2}\mathcal{O},\ldots\}$, meaning that $F(t)$ encodes, in principle, all the information needed to extract $\{b_{n}\}$ [80]. Since the polynomial fit requires accurate estimation of the autocorrelation function, which becomes challenging at large $t$ due to decoherence in the experimental setup, only early time KC estimation is likely to be accessible. These still can capture certain distinguishing characteristics of chaotic dynamics (Fig. 11).
VI Conclusion and Outlook
@@
In this work, we have investigated the emergence of quantum chaos and complexity in the mixed-field Ising model defined on Erdős–Rényi graphs, and studied how the connectivity $\tilde{M}$ governs the onset of chaos. Hamiltonians of similar structure appear naturally in quantum circuits related to optimization tasks, and careful use of such Hamiltonian terms can have a catalytic effect in improving the performance of QAOA and annealing algorithms (Sec. II, also see Ref. [90] ). Using a combination of diagnostics – deep thermalization of the projected ensemble, the partial spectral form factor, and Krylov complexity, we provide a detailed characterization of the localized-to-chaotic-to-integrable crossover driven by changes in connectance. All metrics consistently point to the onset of chaotic behavior at intermediate $\tilde{M}$, even for modest system sizes. With increasing $L$, the chaotic region in the space of $\tilde{M}$ expands, and we expect that in the thermodynamic limit, the model is chaotic everywhere except at $\tilde{M}=0$ and $1$.
Deep thermalization of the projected ensemble reveals a rapid approach of PE to Haar ensemble with system-size at intermediate $\tilde{M}\sim 0.4-0.7$ indicating chaotic dynamics. Away from intermediate $\tilde{M}$, the metrics of deep thermalization, namely the trace distance of PE moments from that of the Haar ensemble, show a smooth crossover from chaotic to integrable/localized features. This is reflected in the time dependence of the trace distances as well as the system size dependence of the asymptotic plateau in the trace distance. The $\tilde{M}\sim 0$ systems have disconnected or weakly connected clusters, which result in local integrals of motion. At $\tilde{M}=1$, i.e., the LMG limit, the system commutes with the total spin operator $S^{2}$ as well as has a large number of permutation symmetries, causing fragmentation of the Hilbert space into fragments of sizes at most $L+1$. Away from this limit, the deletion of a finite density of bonds on the LMG model results in weak mixing of these fragments. Thus, just away from the extreme limits, we expect approximate conservation laws that cause slow relaxation. These also cause non-ergodic eigenstates, which are reflected in the asymptotic trace distance plateau being different from that of the maximally chaotic $\tilde{M}\sim 0.4$. These diagnostics can be scaled up in currently accessible quantum simulators, making our finite-size results a baseline for such an experimental effort. Construction of the projected ensemble, estimation of its higher moments, and behavior of deep thermalization measures have already been demonstrated on quantum simulator platforms with up to $25$ qubits in both neutral-atom arrays [28] and superconducting architectures [31]. The finite-size signatures of deep thermalization reported here, such as the $\tilde{M}-$dependent exponent in the time-dependent decay of the trace distance, are within reach of current systems. With improvements in coherence time, larger system sizes will become accessible on quantum devices, enabling extrapolations toward the thermodynamic limit.
We have shown that the partial spectral form factor also provides a clear characterization of the chaotic dynamics of the random graph Hamiltonian, similar to what has been achieved using the spectral form factor. The intermediate $\tilde{M}$ graphs agree well with the chaos imprints exhibiting a correlation hole, a ramp, and a plateau structure. The localizing Hamiltonian in the small $\tilde{M}$ regime has a pSFF without a correlation-hole structure. Both $\tilde{M}\sim 1$ and $\tilde{M}\sim 0$ systems show larger fluctuations and differ in the mean value of the late-time plateau of the pSFF. Overall, the behavior of the pSFF in the localizing, integrable, and chaotic regimes is qualitatively similar to that of the spectral form factor. While SFF and pSFF can be estimated using randomized measurements, pSFF requires measurements and controlled unitaries only at a small subsystem level [52, 34]. The better experimental scalability makes pSFF a valuable tool for investigating the emergence of chaos in larger systems using quantum simulators. Realizing the full ramp-and-plateau structure of the pSFF does, however, require long coherence times to evolve the system to $t\sim t_{p}$, which scales exponentially with $L$. The correlation hole appears considerably earlier $t_{\mathrm{Th}}$ and serves as a potential discriminator between chaotic and non-chaotic regimes.
The relaxation dynamics studied in Appendix F also show distinct long-time behavior in the different connectance regimes. While the chaotic regime shows anomalous thermalization behavior reminiscent of the quantum Mpemba effect, the integrable regime shows persistent oscillations in the late-time trace distance from the equilibrium density matrix.
Krylov complexity provides the cleanest distinction of dynamical regimes, showing clearly larger values at intermediate connectance consistent with chaos (Fig. 12(c)). However, estimating KC experimentally is challenging at present. Experimentally measurable proxies have been proposed for KC; these can help scale up the study of the emergence of chaos from an information scrambling perspective. An analogous investigation applicable directly to the quantum annealing setting requires analysis of the low-energy sector of the spectrum using probes such as an energy-density-specific Krylov complexity, as the performance of quantum annealing is sensitive to the localization of low-lying eigenstates [20]. Overall, our results provide a comprehensive and experimentally relevant characterization of the connectivity-driven onset of quantum chaos, offering benchmarks and insights for controlling the performance of variational quantum algorithms and reservoir computing [50].
@@ Acknowledgements.
Appendix A Eigenstate Localization and Entanglement
Here, we briefly discuss the eigenstate properties of the Hamiltonian in Eq. 1 with $g=1$ and $h=0.1$ for a $L=10$. We take a single realization of the Hamiltonian at each $\tilde{M}$ and compute the bipartite entanglement entropy of each eigenstate. We quantify the entanglement entropy with Von-Neumann entropy, $R_{1}^{A}=-\sum_{i}\lambda_{i}\log\lambda_{i}$, where $\lambda_{i}$ are the squared singular values of the subsystem reduced density matrix. The result is presented in Fig. 14 for a range of connectivities. The maximum entanglement is bounded by Page value, $R_{1}^{\rm{A,Page}}=(L/2)\log 2-0.5$ and is marked with the black dashed line. The entanglement is smaller than Page value even for the mid-spectrum states at $\tilde{M}$ close to 0. For $\tilde{M}\approx 0.3-0.7$, the entanglement entropy of most of the bulk eigenstates approaches the Page value. As $\tilde{M}\to 1.0$, some fraction of mid-spectrum eigenstates show lower entanglement entropy. At $\tilde{M}=1.0$, signs of fragmentation appear, with eigenstate entanglement values clustering.
Figure 14: The bipartite entanglement entropy for a single instance of the Ising problem on an ER graph of $L=10$ sites with $L_{A}=5$. The black dashed line shows the Page value.
Figure 15: The IPR of product states (a) $|\psi^{(1)}\rangle$ (b) $|\psi^{(2)}\rangle$. All data points are averaged over $100$ graph instances (and initial state in the case of (b))
Next, we examine the localization of simple product states in the eigenstate basis. We measure this quantity with inverse participation ratio (IPR), defined as
$\mathrm{IPR}(\psi)=\sum_{i=1}^{d}|\langle E_{i}|\psi\rangle|^{4},$ (26)
where $E_{i}$ are Hamiltonian eigenstates. IPR is $1$ for perfectly localized state and $1/2^{L}$ for maximally delocalized states. We compute the IPR for two classes of initial product states,
$|\psi^{(1)}\rangle$ $=\frac{1}{\sqrt{2^{L}}}\bigotimes_{i=0}^{L}(|0\rangle+\imath|1\rangle),$ (27)
These states (same as Eq. 11) have 0 energy and are permutation-invariant. Next, we consider a class of states that lack the permutation invariance. Consider each site $i$ is prepared in an equatorial state on the Bloch sphere,
$|\psi^{(2)}\rangle=\bigotimes_{i=1}^{L}|\phi_{i}\rangle,\qquad|\phi_{i}\rangle=\frac{|0\rangle+e^{i\phi_{i}}|1\rangle}{\sqrt{2}},$ (28)
where $|0\rangle$ and $|1\rangle$ are the eigenstates of $Z$. The angles $\{\phi_{i}\}$ are drawn from antipodal pairs to ensure zero energy. Specifically, $L/2$ angles $\alpha_{k}\sim\mathrm{Uniform}[0,2\pi)$ are sampled independently, giving the set
$\{\phi_{i}\}_{i=1}^{L}=\{\alpha_{1},\,\alpha_{1}+\pi,\;\alpha_{2},\,\alpha_{2}+\pi,\;\ldots,\;\alpha_{L/2},\,\alpha_{L/2}+\pi\},$ (29)
randomly assigned across sites. The zero-energy condition follows from the cancellation within each antipodal pair,
$\langle H\rangle=-g\sum_{i=1}^{L}\cos\phi_{i}=-g\sum_{k=1}^{L/2}\!\Big(\cos\alpha_{k}+\cos(\alpha_{k}+\pi)\Big)=0,$ (30)
where the $J_{z}$ and $h_{z}$ terms vanish identically since $\langle\sigma^{z}_{i}\rangle=0$ for all equatorial states. The special case $\alpha_{k}=\pi/2$ for all $k$ recovers the uniform $|+y\rangle^{\otimes L}$ initial state used in the main text in Eq. 11. Fig. 15(a) shows the IPR for $|\psi^{(1)}\rangle$ averaged over $100$ Hamiltonians. The IPR is minimum near $\tilde{M}=0.3$, indicating the state is maximally spread among the eigenstates compared to the case of other $\tilde{M}$. Incidentally, $\Delta^{(k)}$ attains its minimum value around this $\tilde{M}$ in Fig. 6. The IPR goes up (the product state is localized in the eigenbasis of the Hamiltonian) at extremal values of $\tilde{M}$.
For $|\psi^{(2)}\rangle$, we average over both initial state and Hamiltonians at each $\tilde{M}=0.3$. As shown in Fig. 15(b), for larger systems, there is a clear minimum in IPR at $\tilde{M}=0.5$. At very small systems ($L=6$), this distinction gets weak.
Appendix B Additional data on QAOA with random Ising mixer
In Sec. II.2, we provided an illustrative example demonstrating the effect of the additional mixer Hamiltonian on QAOA performance for $L=8$ and $\tilde{M}_{c}=0.4$. Here, we present the QAOA performance for an random Ising problem instance on an ER graph with $L=10$ and $\tilde{M}_{c}=0.3$ to validate the results on larger system. The $H_{d}$ was taken with Pauli-$YY$ coupling (Eq. 8). This $YY$ term can scatter among computational basis states with Hamming distance $2$ compared to just the transverse field mixer which scatters among states with Hamming distance $1$. We demonstrate the numerical results in Fig. 16 for $500$ random instances of $H_{d}$ at each $\tilde{M}$. Here, we first define a baseline for performance comparison. We consider the best $E_{\text{QAOA}}$ obtained with $\tilde{M}\in\{0,1\}$ as $E^{\text{ref}}_{\text{QAOA}}$. The red line denotes the fraction of instances where the energy obtained with QAOA protocol is lower compared to $E^{\text{ref}}_{\text{QAOA}}$. The fraction of instances surpassing the baseline energy shows two distinct trends: at fixed $p$, it is maximized at intermediate $\tilde{M}$, confirming that the chaotic regime provides a favorable landscape for QAOA optimization; additionally, at any fixed $\tilde{M}$, this fraction grows with $p$, indicating that deeper circuits with more applications of the chaotic driver Hamiltonian progressively improve the ability to explore the energy landscape and find lower-energy solutions. The latter trend is consistent with the expectation that repeated application of a chaotic driver generates more complex quantum superpositions, allowing QAOA to access a larger portion of the Hilbert space with increasing circuit depth. This trend empirically holds for circuits with up to $p=3$ layers that we have studied in this work. However, very deep quantum circuits suffer from barren plateaus where the cost function gradient vanishes exponentially with system size, making classical parameter optimization intractable.
Figure 16: The approximation ratio for an Ising problem on an ER graph of $L=10$ sites with $\tilde{M}_{c}=0.3$. The red line shows the fraction of QAOA instances where the corresponding connectivity performs better than the baseline.
As discussed in Sec . II, the improvement in QAOA performance observed with the $H^{YY}$ driver could stem from the chaotic nature of $H_{\rm d}$, and the presence of terms that scatter among computational basis states more efficiently. We consider an alternative driver Hamiltonian with Pauli-$ZZ$ coupling,
$H_{\rm d}$ $=-\frac{JL}{M}\sum_{i,j}A_{ij}Z_{i}Z_{j}-g\sum_{i}X_{i},$ (31)
which is diagonal in the computational basis and therefore induces no scattering among basis states. Taking the same problem instance $H_{\rm c}$ as in Fig. 4(a),(c) (Problem 1), we present the distribution of QAOA performance for $p=2,3$ and $4$ in Fig. 17. The improvement in performance with intermediate $\tilde{M}$ persists with the $ZZ$ driver, despite the absence of any Hamming distance $2$ scattering. This establishes that the chaotic structure of $H_{\rm d}$, rather than the specific scattering properties of the mixer, is the primary driver of the observed improvement in QAOA.
Figure 17: The approximation ratio for an Ising problem on an ER graph of $L=8$ sites with $\tilde{M}_{c}=0.4$ (same $H_{\rm c}$ as Problem 1 in Fig. II.2). The red line shows the fraction of QAOA instances where the corresponding connectivity performs better than the baseline.
Appendix C PE from a Permutation-Symmetry-Breaking Initial State
Figure 18: (a)-(c) Asymptotic ($t=10^{9}J^{-1}$) trace distance of first 3 moments of PE from Haar ensemble as a function of $L_{B}$. The asymptotic trace distance gets suppressed with increasing $L_{B}$ for all $\tilde{M}$. Error bars represent the standard error of the mean across realizations at a fixed $t$. (d)-(f) The same data plotted to show the the scaling of $\overline{\Delta^{(k)}}(t\to\infty)$ with $L$ across all $\tilde{M}$ for first $3$ moments of PE. The error bars are not shown when they are comparable to the marker size. Each data point shown represents an average over 500 graph realizations and random direct product initial states (Eq. 28), calculated for a subsystem of size $L_{A}=2$ with Hamiltonian parameters (Eq. 1) $g=1.0$ and $h=0.1$. The black dashed line shows the trace distance of an empirical ensemble of $500$ random states (sampled from Haar random states of system size $L$) from the corresponding moments of the Haar ensemble.
The initial state is chosen to be a product state $\psi^{(2)}$ (Eq. 28 with zero energy expectation value, $\langle\psi_{0}|H|\psi_{0}\rangle=0$ and without permutation symmetry.
Figure 18 shows the scaling of trace distance between the first $3$ moments of PE and Haar ensemble across connectivity and system size. The initial state is taken as a random direct product state (Eq. 28) with energy expectation value 0. The data is presented for $500$ random initial states and ER graphs. The qualitative behavior is similar to Fig. 6.
Appendix D Lanczos algorithm